A ball starts from origin at time t=0 secs with an initial velocity in x-component 30m/s and a y-component is -25m/s. So, the ball moves in xy-plane with acceleration 4.2. Find the total velocity at time t=3.2 seconds
1
Expert's answer
2019-12-16T11:01:37-0500
"v_0=\sqrt{30^2+(-25)^2}=39.05\frac{m}{s}"

The magnitude:

"v=v_0+at=39.05+4.2(3.2)=52.5\frac{m}{s}"

The direction is

"\theta=\arctan{\frac{25}{30}}=39.8\degree\ below\ x-axis"